If x + y = 1.3 and x - y = 0.9, what is y?
I've tried to work this out, I know that 1.3-x=y, and 0.9-y=x, but I am lost as to where to go from there. Thank you for any assistance as to how to go about solving this.

Answer by ewatrrr(24785)

 
Hi
This System of EQuations lends to solving by first Eliminating the y
x + y = 1.3
x - y = 0.9
2x = 2.2
x = 1.1 and y = .2 ||As You listed y = 1.3-x